OpenStudy (anonymous):

College Algebra: Inequalities In Two Variables! I need to sketch the graph of the nonlinear inequality. Sadly, I'm not good at reading and understanding how to graph these. The math problem is: y>-x^2

OpenStudy (anonymous):

Okay, so draw the line y=-x^2 and we get this: Try a point above the line, say (0, 5). Does this satisfy your inequality? y>-x^2? Try a point below, say (0, -5), does this satisfy your inequality?

OpenStudy (anonymous):

From these two points you will spot for which side of the line the inequality is true for :)

OpenStudy (anonymous):

How do you graph y>-x^2? Like how do you read it in order to graph it?

OpenStudy (anonymous):

I haven't, I plotted y=-x^2. I know what y=x^2 looks like, and I know that y=-x^2 is hte same but reflected in the x axis.

OpenStudy (anonymous):

I understand what you're saying though and I see that I need to shade everything around the upside down U. Oh, I'm the kind of person who doesn't know what y=-x^2 looks like on a graph so I wouldn't know how to graph it. I have a Chapter Test tomorrow and I'm trying to learn a few things in case they pop up.

OpenStudy (anonymous):

If you don't know what a graph looks like, try creating a table of values then plotting it: